Write the slope intercept form of the equation of the line that passes through the point (7, 14), and is parallel
to - 2x + 3y = 4 .

Answer by mananth(16946)

-2 x + 3 y = 4
Find the slope of this line
make y the subject
3 y = 2 x + 4
Divide by 3
y = 2/3 x + 1.33
Compare this equation with y=mx+b
slope m = 2/3
The slope of a line parallel to the above line will be the same
The slope of the required line will be 2/3
m= 2/3 ,point ( 7 , 14 )
Find b by plugging the values of m & the point in
y=mx+b
14 = 4 2/ 3 + b
b= 28/3
m= 2/3
Plug value of the slope and b in y = mx +b
The required equation is y =(2/3)x+28/ 3